Travis is 21 years older than his sister Michelle. In 9 years, he will be twice as old as Michelle. How old is each of them now?
Found 2 solutions by stanbon, greenestamps:
Answer by stanbon(75887)

Travis is 21 years older than his sister Michelle. In 9 years, he will be twice as old as Michelle. How old is each of them now?
------------------
Equations:
T = M + 21
T+9 = 2(M+9)
--------
Substitute for "T" and solve for "M":
M+21+9 = 2M + 18
-----
M + 30 = 2M+18
M = 12 (Michelle's age now)
-------------
T = M + 21 = 33 (Travis's age now)
-------------
Cheers,
Stan H.
------------

Answer by greenestamps(13200)



You can solve this without formal algebra by using logical reasoning.

Travis will always be 21 years older than is sister.
He will be twice as old as his sister when she is 21 and he is 42.
The problem says that will happen 9 years from now, so his age now is 42-9 = 33.